OpenStudy (thatonegirl_):

Can someone plz help? How do I graph y=1-πcos1/4x???

OpenStudy (babynini):

https://www.desmos.com/calculator :)

OpenStudy (thatonegirl_):

Yeah but without a calculator?

OpenStudy (babynini):

Er...I'm not for sure. Sorry! :/

OpenStudy (thatonegirl_):

Its okay, thx anyways :)

OpenStudy (loser66):

First off, \(y= 1-\pi cos\dfrac{1}{4x}\) or \(y = 1-\pi cos \dfrac{x}{4}\)??

OpenStudy (thatonegirl_):

\[y=1-\pi \cos \frac{ 1 }{ 4 }x\]

OpenStudy (loser66):

ok. Make a table like this